4. (HHM Problem 2.1 #2) Assume that a vowel is one of the five letters A, E, I, O, or U. (a) How many eleven-letter sequences from the alphabet contain exactly three vowels? (b) How many of these have at least one repeated letter?
Answer #1

(a)

Number of ways of selecting 3 positions for vowels out of 11 is

C(11,3) = 165

Now these three vowels can be filled in 5*5*5 = 125 ways

And rest 11-3 = 8 letters can be filled by 26-5=21 consonanats in 21^{8} ways.

So possible number of 11 letter sequence is

165. 125.218

(b)

Number of 11 letter sequence : 26^{11}

Number of 11 letter sequence with all letters different is : 26\cdot 25\cdot 24...\cdot 16

Number of 11 letter sequence with at least one repeated letter is: 26^{11}-\left ( 26\cdot 25\cdot 24...\cdot 16 \right )